# Automated Patient Positioning with Learned 3D Hand Gestures

### v1

@@ 2 Related Work
Deep learning-based approaches are the current state-of-the-art. Some methods [1, 26] use convolutional neural networks to classify the gestures directly. However, these method do not generalize well in the wild. 3D hand pose estimation methods [2, 7] first regress the 3D coordinates of hand keypoints, which can further be used for hand gesture recognition. However, the processing time of these methods is the bottleneck for real-time applications.
@@ 4.3.1 Ablation Study
Heatmap “okay” “non-okay” AUC
@@ 4.3.2 Comparison with State-of-the-Art
We further conduct a comparison of the run-time complexity between our method and MediaPipe, as illustrated in Table 1. We can see that MediaPipe achieves the highest FPS rate. Nevertheless, our approach demonstrates comparable run-time complexity (14.4 vs. 17.4 FPS), while achieving significantly better results in terms of AUC (0.948 vs. 0.754). This indicates the superior performance and efficacy of our method.
@@ 4.3.3 Evaluation on the HaGRID dataset
To further evaluate our method, we conduct experiments on the HaGRID [10] dataset. First, we crop the hand images from the HaGRID dataset. Second, we train our two-layer MLP gesture classifier on the cropped training dataset using the same settings as HAGRID [10] (i.e., SGD optimizer with a learning rate of 0.1 and a batch size of 128) for 10 epochs. As shown in Table 4, our gesture classifier archives a higher F1-score than HaGRID (98.3 vs. 99.7). At last, our whole pipeline: hand detection, 2D landmark detection, and dual-modality 3D hand pose and gesture classification, achieves higher mAP (79.1 vs. 85.0) for hand detection and gesture recognition compared with HaGRID.
Table 4: Comparison of gesture detection and classification on the HaGRID [10] dataset.
@@ 5 Discussion and Conclusion
We proposed a novel workflow and analytics pipeline to automatically detect and recognize gestures from technicians in medical scans for automated patient positioning. This workflow offers numerous benefits, including increased efficiency, improved patient comfort, more consistent and accurate positioning, and potentially improved image quality. We validate the workflow and analytics pipeline on data captured by RGB-Depth camera (only RGB images used for 3D hand gesture recognition) in a realistic clinical environment extensively, demonstrating promising reliability and positioning accuracy.

### v2

@@ 2 Related Work
Deep learning-based approaches are the current state-of-the-art. Some methods [1, 27] use convolutional neural networks to classify the gestures directly. However, these methods do not generalize well in the wild. 3D reconstruction and pose estimation methods [2, 8, 6] can regress the 3D coordinates of keypoints, which can further be used for hand gesture recognition. However, the processing time of these methods is the bottleneck for real-time applications.
@@ 4.3.1 Ablation Study
Heatmap “okay" “non-okay" AUC
@@ 4.3.2 Comparison with State-of-the-Art
We further conduct a comparison of the run-time complexity between our method and MediaPipe, as illustrated in Table 1. We can see that MediaPipe achieves the highest FPS rate. Nevertheless, our approach demonstrates comparable run-time complexity (14.4 vs. 17.4 FPS), while achieving significantly better results in terms of AUC (0.948 vs. 0.754). This indicates the superior performance and efficacy of our method. Note that, all of our testing, including the FPS measurements, is conducted directly on an NVIDIA Jetson Xavier NX edge device.
@@ 4.3.3 Evaluation on the HaGRID dataset
To further evaluate our method, we conduct experiments on the HaGRID [11] dataset which includes 18 gesture classes in total. First, we crop the hand images from the HaGRID dataset. Second, we train our two-layer MLP gesture classifier on the cropped training dataset using the same settings as HAGRID [11] (i.e., SGD optimizer with a learning rate of 0.1 and a batch size of 128) for 10 epochs. As shown in Table 4, our gesture classifier archives a higher F1-score than HaGRID (98.3 vs. 99.7). At last, our whole pipeline: hand detection, 2D landmark detection, and dual-modality 3D hand pose and gesture classification, achieves higher mAP (79.1 vs. 85.0) for hand detection and gesture recognition compared with HaGRID.
Table 4: Comparison of gesture detection and classification on the HaGRID [11] dataset with 18 gesture classes.
@@ 5 Discussion and Conclusion
In this paper, we introduce a novel workflow and analytics pipeline for automatically detecting and recognizing gestures from technicians during medical scans, to automate patient positioning. Our system uses a ceiling-mounted RGB-Depth camera, relying solely on the RGB component for 3D hand gesture recognition, and is extensively tested in a realistic clinical environment. Additionally, our 3D hand gesture recognition method is evaluated on the HaGRID public dataset, achieving state-of-the-art performance.
Future work will include clinical trials to quantify the potential benefits and compare our system’s performance with traditional manual positioning methods, allowing us to validate our method’s impact on efficiency, accuracy, and patient comfort.